OpenStudy (anonymous):

Given the function g(x) = −2x + 4, find g(−5). −6 4.5 6 14

OpenStudy (anonymous):

substitute -5 in to x g(x) = −2x + 4, g(-5)= -2(-5)+4 g(-5)= 10+4 g(-5)= ? Hope this helps

OpenStudy (anonymous):

14

OpenStudy (anonymous):

correct